What is the term for the pressure that creates water flow between two points?
Answer #1

Bernoulli's principle in fluid dynamics gives the pressure that creates water flow between two points.

Bernoulli's principle can be derived from the principle of conservation of energy.

Consider a stream line of steady flow of an ideal fluid along a stream tube at points A and B.

Apply Bernoulli's principle at A and B,

\frac{P_{A}}{\rho g}+\frac{V_{A}^{2}}{2g}+Z_{A}=\frac{P_{B}}{\rho g}+\frac{V_{B}^{2}}{2g}+Z_{B}

Pressure difference between two points:

\frac{P_{A}-P_{B}}{\rho g}=(\frac{V_{B}^{2}-V_{A}^{2}}{2g})+(Z_{B}-Z_{A})